Answer the following genetics problems. Show your work!

A parent plant with alleles for colors of flower, seed, and leaf had a genotype of AaBbCc. It was crossed with the flower of genotype AaBbCc. What fraction of offspring are predicted to be homozygous recessives for at least two of the three characters? (Assume independent assortment of all gene pairs)
Answer #1

Answer:

AaBbCc x aaBbCc –Parents

Aa x Aa = AA (1/4), Aa (1/2) & aa (1/4)

Bb x Bb = BB (1/4), Bb (1/2) & bb (1/4)

Cc x Cc = CC (1/4), Cc (1/2) & cc (1/4)

Probability of A_bbcc = ¾ * ¼ * ¼ = 3/64

Probability of aaB_cc = 1/4 * 3/4 * ¼ = 3/64

Probability of aabbC_ = 1/4 * ¼ * 3/4 = 3/64

Probability of aabbcc = ¼*1/4*1/4 = 1/64

fraction of offspring are predicted to be homozygous recessives for at least two of the three characters = 3/64 + 3/64 + 3/64 + 1/64 = 10/64